# Gradient Descent with Large Step Sizes: Chaos and Fractal Convergence Region

### v1

@@ 3.2 Regularization Induces Fractal Convergence Boundary
Let $(u_{t},v_{t})_{t\geq 0}$ denote the gradient descent trajectory in problem (4) with $\lambda\geq 0$. Let $(z_{t},w_{t})=T(u_{t},v_{t})$. There exists a planar map $F\colon\mathbb{R}^{2}\to\mathbb{R}^{2}$ that only depends on $\eta,\lambda,y$ such that $(z_{t+1},w_{t+1})=F(z_{t},w_{t})$ holds for all $t\geq 0$. In particular, $(u_{t},v_{t})$ converges to $\mathcal{M}$ if and only if $(z_{t},w_{t})$ converges to $\{z=y\}$, and it converges to $(\mathbf{0},\mathbf{0})$ if and only if $(z_{t},w_{t})$ converges to $(0,0)$.
@@
[…]
@@ A.1 Chaotic dynamical systems
A common definition of chaotic dynamical systems is by Devaney and Eckmann, (1987). A dynamical system $F\colon X\to X$, where $X$ is the state space, is chaotic if: (i) it is sensitive to initialization, (ii) it is topological transitive, and (iii) periodic points are dense in $X$. Here, sensitivity to initialization requires that there exists $\delta>0$ such that for any $x\in X$ and any neighborhood of $x$, there exists $N>0$ and $y$ in the neighborhood such that $\operatorname{dist}(F^{N}(x),F^{N}(y))>\delta$. This is weaker than the property we presented in Theorem 1 and Theorem 3: when the converged points of two trajectories are different, the trajectories must differ by a positive difference at some time $N$, but not vice versa. In Proposition 14, we show that the boundary $\partial\mathcal{D}_{\eta}^{\prime}$, as defined in Theorem 1, is invariant under gradient descent. In Proposition 18, we show that when restricted to $\partial\mathcal{D}_{\eta}^{\prime}$, the gradient descent system is semi-conjugate to a one-dimensional system that is precisely Devaney chaotic. However, we show in Proposition 19 that the original gradient descent system is not Devaney-chaotic when $d\geq 2$ as it fails to be topological transitive.
@@ Appendix D Quotient dynamics of gradient descent
$T\colon\mathbb{R}^{2d}\to\mathbb{R}^{2},\ T(u,v)=(u^{\top}v-y,\left\lVert u\right\rVert^{2}+\left\lVert v\right\rVert^{2}).$
@@ D.1 Quotient dynamical system
We have that $(z_{t+1},w_{t+1})=f(z_{t},w_{t})$ holds for all $t\geq 0$. In particular, $(u_{t},v_{t})$ converges to $\left\{u^{\top}v=y\right\}$ if and only if $(z_{t},w_{t})$ converges to $\{z=0\}$, and it converges to $(\mathbf{0},\mathbf{0})$ if and only if $(z_{t},w_{t})$ converges to (y,0)(-y,0).
@@
Note that, the loss function solely depends on $u^{\top}v-y$. Thus $(u_{t},v_{t})$ converges to $\left\{u^{\top}v=y\right\}$ if and only if $(z_{t},w_{t})$ converges to $\{z=0\}$. Also, note that, T(u,v)=(y,0)T(u,v)=(-y,0) if and only if $u=v=\boldsymbol{0}$. Thus $(u_{t},v_{t})$ converges to $(\boldsymbol{0},\boldsymbol{0})$ if and only if $(z_{t},w_{t})$ converges to (y,0)(-y,0). This completes the proof. ∎
@@ D.2 Properties of the quotient dynamics
and $\left\{Q(z,w)>r\right\}=\Omega\setminus\left\{Q(z,w)=r\right\}$.
@@
Therefore, the line $16\mu z-r^{2}+2rw=0$ intersect with $\partial\Omega$ at two points, whose $w$ coordinates are $\frac{r\pm 4\mu}{2}$. Since $r>4|\mu|$, we have that $\frac{r\pm 4\mu}{2}<r$ always holds. This implies that, for all $(z,w)\in\Omega\cap\left\{-16\mu z-r^{2}+2rw<0\right\}$, we have rw>0r-w>0. Thus, we have that for $(z,w)\in\Omega$ and $r>4|\mu|$,
@@
The above clearly hold when << is changed to >>. This completes the proof. ∎
We identifies three invariant sets of the quotient system $F$. Recall that, a set $S\subset\Omega$ is said to be an invariant set under $F$ if $F(S)\subset S$.
@@
When $\nu>0$, we have that, for $(z,w)\in\Omega$, $Q(F(z,w))\leq Q(z,w)$ if and only if $(z,w)$ lies in the set
$\{z^{2}\leq-\nu^{2}+2\nu\}\cup\left\{w<-\frac{\mu(z^{2}-2\nu+\nu^{2})}{z(\nu-1)}-\frac{4z^{2}(\nu-1)}{\nu^{2}-2\nu+z^{2}},z^{2}>-\nu^{2}+2\nu\right\}.$
In particular, $Q(z,w)\geq F(Q(z,w))$ if $Q(z,w)<8-4\nu$.
@@
Let $(z^{\prime},w^{\prime})=F(z,w)$. Assume that $\mu>0$. The case of $\mu<0$ can be proved via a analogous procedure. Let $r=Q(z,w)$ and let $s=w-\sqrt{w^{2}-16\mu z}$. When $r=4|\mu|$, we have that $Q(F(z,w))=Q(z,w)$ always holds, by Lemma 11. Consider $r>4|\mu|$. Using Lemma 10, we have that the sign of $Q(z^{\prime},w^{\prime})-Q(z,w)$ is the same as that of the inner product between the vector pointing from $(z,w)$ to $(z^{\prime},w^{\prime})$ and the normal vector $(-8\mu,Q(z,w))$ of the line $16\mu z-Q(z,w)^{2}+2Q(z,w)w=0$, which is given by
@@
$\{2\mu(z^{2}+\nu^{2}-2\nu)-(1-\nu)z(w-\sqrt{w^{2}-16\mu z})<\leq 0\}$
$=$ $\{z^{2}\leq-\nu^{2}+2\nu\}\cup\left\{w<-\frac{\mu(z^{2}-2\nu+\nu^{2})}{z(\nu-1)}-\frac{4z^{2}(\nu-1)}{\nu^{2}-2\nu+z^{2}},z^{2}>-\nu^{2}+2\nu\right\}.$
@@
When $\mu=0$, $Q(z,w)=2w$. We have that
@@
Notice that $p(-\infty)=$-$\infty$ and $p(-1)=(\nu-1)^{2}(w_{0}-2(z_{0}+\mu))>0$. Hence, $p$ has at least one root in $(-\infty,-1)$. By Lemma 11, in particular, by (9), we have that, when viewing $F$ as a map on $\mathbb{R}^{2}$:
@@
We now show that $F$ maps $\partial A_{0}$ bijectively to $\partial\Omega$. Since $A_{0}\subset\mathrm{cl}(A_{0})$, we have that $F(A_{0})=\Omega^{o}\subset F(\mathrm{cl}(A_{0}))$. Since $\mathrm{cl}(A_{0})$ is compact, therefore, $F|_{\mathrm{cl}(A_{0})}\colon\mathrm{cl}(A_{0})\to\Omega$ is proper, and hence is closed (see, e.g., Theorem 4.95. Lee,, 2000). Therefore, $F(\mathrm{cl}(A_{0}))$ is a closed set that contains $\Omega^{o}$. Hence, $\mathrm{cl}(\Omega^{o})=\Omega\subset F(\mathrm{cl}(A_{0}))$. Since $F(A_{0})=\Omega^{o}$, we have $\partial\Omega\subset F(\partial A_{0})$, which means $F|_{\partial A_{0}}$ is onto $\partial\Omega$. By Lemma 11, we know $F$ maps $\left\{z=\nu-1\right\}$ to $\left\{w=2(z+\mu)\right\}$ and maps $\left\{w=-2(z+\mu)\right\}$ to itself. When $z=\nu-1$, the $w$-update is given by
@@
which is linear in $w$. Therefore, $F|_{z=1-\nu}$ must be an injection. When $w=-2(z+\mu)$, the $w$-update is given by
@@
As a function $w$, $w^{\prime}$ have two critical points, $w=2(1-\mu-\nu)$ and $w=\frac{2}{3}(1-\mu-\nu)$. Notice that $z=\nu-1$ intersects $\partial\Omega$ at $(\nu-1,2(1-\mu-\nu))$. Then when $(z,w)\in\mathrm{cl}(A_{0})$, we have $w\geq 2(1-\mu-\nu)$. Since $1-\mu-\nu>0$, we have that the above $w^{\prime}$ is monotonic with $w$. Therefore, $F|_{\mathrm{cl}(A_{0})}$ is an injection. It follows that, $F|_{\partial A_{0}}$ is a bijection to $\partial\Omega$ and $F|_{\mathrm{cl}(A_{0})}$ is a bijection to $\Omega$. Note, $F|_{\mathrm{cl}(A_{0})}$ is proper by Proposition 9 and by $\mathrm{cl}(A_{0})$ is closed. Hence, $F|_{\mathrm{cl}(A_{0})}$ is a closed map (see, e.g., Theorem 4.95. Lee,, 2000), which means its inverse is continuous. Hence, $F|_{\mathrm{cl}(A_{0})}$ is a homeomorphism. The proof for $A_{2}$ is similar and thus is omitted.
@@
$6-4\nu-\frac{1-3z^{2}-2z\mu-2\nu+\nu^{2}}{1-\nu}>0$
$\Leftrightarrow\$$3z^{2}+2\mu z+(3\nu^{2}-8\nu+5)>0.$
For this parabola, we have
@@
So the parabola is always above $y=0$. Therefore, we have that
@@
For this new parabola, we have its discriminant is negative if
@@ Appendix E Proofs for Section 3
In this section, we present the proofs of our main results. The key idea is to first analyze the quotient dynamical system introduced in Appendix D, and then translate the conclusions back to the original system.
@@ E.1.1 Preliminary results
As discussed in Appendix D, the gradient descent dynamics are captured by the following system $F$:
@@
where $\mu=y\eta$ denote the parameter of system $F$. Given $\eta>0$, the state space of $F$ is
@@
In the following several results, we characterize the long term behavior of orbits of $F$. We will use the terms trajectory and orbit interchangeably. We say an orbit $\{x_{k}\}$ converges to a set $S$ if $d(x_{k},S)\to 0$, where $d(x,S)=\inf_{y\in S}d(x,y)$. Unless stated otherwise, we use $(z^{\prime},w^{\prime})$ to denote $F(z,w)$.
@@
  • If $w_{0}>\mu z_{0}+4$, the orbit either diverges, in the sense that $w_{k}\rightarrow\infty$ and $z_{k}\not\to 0$, or converges to $\{z=0\}$ in finite steps.
@@
Consider function $\Delta Q(z,w)=Q(F(z,w))-Q(z,w)$. When $w_{0}<\mu z_{0}+4$, by Lemma 12, we have $Q(z_{k+1},w_{k+1})\leq Q(z_{k},w_{k})$ for all $k\geq 0$. Hence, the trajectory stays in the region $\{w<\mu z+4\}$. Since $Q(z_{k},w_{k})$ is bounded from below, it converges to some finite value and $\Delta Q(z_{k},w_{k})$ converges to zero. Meanwhile, from Lemma 12,
@@
If $(z_{k},w_{k})$ does not converges to $Z$, there exists $\varepsilon_{0}$ such that for any $K$ there exists $k>K$ such that $d((z_{k},w_{k}),Z)\geq\varepsilon_{0}$. Note the function $\Delta Q$ is uniformly continuous on $\{w<\mu z+4\}\subset\left\{w\leq\mu z+4\right\}$ where the latter is compact. Hence, $d((z_{k},w_{k}),Z)\geq\varepsilon_{0}$ implies that $\Delta Q(z_{k},w_{k})>\delta$ for some $\delta>0$, which contradicts with the fact that $\Delta Q\rightarrow 0$. Therefore, $(z_{k},w_{k})$ converges to
@@
When $w_{0}>\mu z_{0}+4$, similarly, we have that $Q(z_{k+1},w_{k+1})\geq Q(z_{k},w_{k})$. Hence, $(z_{k},w_{k})$ stays in the region $\{w>\mu z+4\}$ for all $k\geq 0$. Hence, $Q(z_{k},w_{k})$ either diverges to infinity or converges to a finite value. If it diverges, $w_{k}$ must also diverge, since the function $Q$ confined on $\{w\leq\bar{w}\}$ for any fixed $\bar{w}$, is continuous and hence upper bounded. Meanwhile, $z_{k}$ can not converge to zero, since points on $\{z=0\}$ are fixed points. If $Q(z_{k},w_{k})$ converges to some finite value $C$, then $\Delta Q$ converges to zero and the trajectory must remain within the region $\{w\leq C^{\prime}\}$ for some $C^{\prime}>0$. Since $\Delta Q$ is uniformly continuous on $\{w\leq 2C^{\prime}\}$, we have $(z_{k},w_{k})$ converges to $Z$, using similar arguments as before. Hence, $(z_{k},w_{k})$ can only converges to $Z\cap\left\{w>\mu z+4\right\}=\left\{z=0\right\}$. Now assume the convergence is in infinite steps, i.e., $|z_{k}|\neq 0$ for all $k\in\mathbb{N}$. Then the sequence $|z_{k+1}/z_{k}|$ is well defined and converges to one. Notice that we have
$|\frac{z_{k+1}}{z_{k}}|=|z_{k}^{2}+\mu z_{k}-w_{k}+1|\geq|z_{k}^{2}+\mu z_{k}|+|w_{k}-1|.$
Since $|z_{k}^{2}+\mu z_{k}|\to 0$, the lower bound is dominated by $|w_{k}-1|$ which is strictly greater than one since $\{w>\mu z+4\}\cap\{z=0\}=\{(0,w)\colon w\geq 4\}$. This contradicts with the fact that $|z_{k+1}/z_{k}|$ converges to one and hence the convergence must be in finite steps.
Finally, the result for the case $w_{0}=\mu z_{0}+4$ directly comes from Lemma 12. With this, we conclude the proof. ∎
@@
First, consider $|\mu|<1$. By Proposition 14, when $Q(z,w)>8$, initializations either converge to $\{z=0\}$ in finite steps or diverge. Notice that converging within finite steps means the initialization lies in the set

### v2

@@ 3.2 Regularization Induces Fractal Convergence Boundary
Let $(u_{t},v_{t})_{t\geq 0}$ denote the gradient descent trajectory in problem (4) with $\lambda\geq 0$. Let $(z_{t},w_{t})=T(u_{t},v_{t})$. There exists a planar map $F\colon\mathbb{R}^{2}\to\mathbb{R}^{2}$ that only depends on $\eta,\lambda,y$ such that $(z_{t+1},w_{t+1})=F(z_{t},w_{t})$ holds for all $t\geq 0$. In particular, $(u_{t},v_{t})$ converges to $\mathcal{M}$ if and only if $(z_{t},w_{t})$ converges to $T(\mathcal{M})$, and it converges to $(\mathbf{0},\mathbf{0})$ if and only if $(z_{t},w_{t})$ converges to $(0,0)$.
@@
Theorem 3 characterizes the geometry of the convergence boundary for problem (4). Specifically, $\partial\mathcal{D}_{\eta}^{\prime\prime}$ can be understood as a fiber-bundle-like object [^2] It is not a rigorous fiber bundle since the fiber $T^{-1}(z,w)$ might degenerate when $(z,w)$ is a singular value.: fibers of $T$, which are generically smooth and bounded manifolds, are attached to a self-similar set $T(\partial\mathcal{D}_{\eta}^{\prime\prime})$ with degree three (see the left panel of Figure 3). The case of $d=1$, i.e., $u,v\in\mathbb{R}$, is illustrated in the middle panel of Figure 3: in this case, the fibers are generically a four-point set and hence, $\partial\mathcal{D}_{\eta}^{\prime\prime}$ simply consists of four copies of $T(\partial\mathcal{D}_{\eta}^{\prime\prime})$. To quantify the fractality, we estimate the box-counting dimension of $T(\partial\mathcal{D}_{\eta}^{\prime\prime})$, which turns out to be $1.249$ (see the right panel of Figure 3). This non-integer dimension implies that $T(\partial\mathcal{D}_{\eta}^{\prime\prime})$ is essentially more complex than any smooth curve, yet fails to occupy any planar area. The fractal boundary marks unpredictability in convergence: when the initial point lies near the boundary, it is practically impossible to determine whether it is inside or outside the convergence region. This unpredictability is also quantified by the box-counting dimension, as explained in Appendix B.2.
@@ A.1 Chaotic dynamical systems
A common definition of chaotic dynamical systems is by Devaney and Eckmann, (1987). A dynamical system $F\colon X\to X$, where $X$ is the state space, is chaotic if: (i) it is sensitive to initialization, (ii) it is topological transitive, and (iii) periodic points are dense in $X$. Here, sensitivity to initialization requires that there exists $\delta>0$ such that for any $x\in X$ and any neighborhood of $x$, there exists $N>0$ and $y$ in the neighborhood such that $\operatorname{dist}(F^{N}(x),F^{N}(y))>\delta$. This is weaker than the property we presented in Theorem 1 and Theorem 3: when the converged points of two trajectories are different, the trajectories must differ by a positive difference at some time $N$, but not vice versa. In Proposition 14, we show that the boundary $\partial\mathcal{D}_{\eta}^{\prime}$, as defined in Theorem 1, is invariant under gradient descent. In Proposition 18, we show that when restricted to $\partial\mathcal{D}_{\eta}^{\prime}$, the gradient descent system is semi-conjugate to a one-dimensional system that is precisely Devaney chaotic. However, we show in Proposition 19 that the original gradient descent system is not Devaney chaotic when $d\geq 2$ as it fails to be topological transitive.
@@ Appendix D Quotient dynamics of gradient descent
$T\colon\mathbb{R}^{2d}\to\mathbb{R}^{2},\ T(u,v)=(u^{\top}v-y,\left\lVert u\right\rVert_{2}^{2}+\left\lVert v\right\rVert_{2}^{2}).$
@@ D.1 Quotient dynamical system
We have that $(z_{t+1},w_{t+1})=f(z_{t},w_{t})$ holds for all $t\geq 0$. In particular, $(u_{t},v_{t})$ converges to $\mathcal{M}$ if and only if $(z_{t},w_{t})$ converges to $T(\mathcal{M})$, and it converges to $(\mathbf{0},\mathbf{0})$ if and only if $(z_{t},w_{t})$ converges to (y,0)(-y,0).
@@
Note that, the loss function solely depends on $u^{\top}v-y$ and $\left\lVert u\right\rVert_{2}^{2}+\left\lVert v\right\rVert_{2}^{2}$. Thus $(u_{t},v_{t})$ converges to $\mathcal{M}$ if and only if $(z_{t},w_{t})$ converges to $T(\mathcal{M})$. Also, note that, T(u,v)=(y,0)T(u,v)=(-y,0) if and only if $u=v=\boldsymbol{0}$. Thus $(u_{t},v_{t})$ converges to $(\boldsymbol{0},\boldsymbol{0})$ if and only if $(z_{t},w_{t})$ converges to (y,0)(-y,0). This completes the proof. ∎
@@ D.2 Properties of the quotient dynamics
and $\left\{Q(z,w)>4|\mu|\right\}=\Omega\setminus\left\{Q(z,w)=4|\mu|\right\}$.
@@
Therefore, the line $16\mu z-r^{2}+2rw=0$ intersect $\partial\Omega$ at two points, whose $w$ coordinates are $\frac{r\pm 4\mu}{2}$. Since $r>4|\mu|$, we have that $\frac{r\pm 4\mu}{2}<r$ always holds. This implies that, for all $(z,w)\in\Omega\cap\left\{-16\mu z-r^{2}+2rw\leq 0\right\}$, we have rw>0r-w>0. Thus, we have that for $(z,w)\in\Omega$ and $r>4|\mu|$,
@@
The above clearly holds when << is changed to == or >>. This completes the proof. ∎
We identifies three invariant sets of the quotient system $F$. A set $S\subset\Omega$ is said to be an invariant set under $F$ if $F(S)\subset S$.
@@
When $\nu>0$, we have that, for $(z,w)\in\Omega$, (i) $Q(F(z,w))\leq Q(z,w)$ if and only if $(z,w)$ lies in the set
$\{z^{2}\leq-\nu^{2}+2\nu\}\cup\left\{w<-\frac{\mu(z^{2}-2\nu+\nu^{2})}{z(\nu-1)}-\frac{4z^{2}(\nu-1)}{\nu^{2}-2\nu+z^{2}},z^{2}>-\nu^{2}+2\nu\right\},$
which contains $\{Q(z,w)<8-4\nu\}$; and (ii) $Q(F(z,w))=Q(z,w)$ if and only if $(z,w)$ lies in the set
$Z\triangleq\left\{Q(z,w)=4|\mu|\right\}\cup\left\{w=-\frac{\mu(z^{2}-2\nu+\nu^{2})}{z(\nu-1)}-\frac{4z^{2}(\nu-1)}{\nu^{2}-2\nu+z^{2}},z^{2}>-\nu^{2}+2\nu\right\}.$
@@
Let $(z^{\prime},w^{\prime})=F(z,w)$. Assume that $\mu>0$. Note the case of $\mu<0$ can be proved via an analogous procedure. Let $r=Q(z,w)$ and let $s=w-\sqrt{w^{2}-16\mu z}$. When $r=4|\mu|$, we have that $Q(F(z,w))=Q(z,w)$ always holds, by Lemma 11. Consider $r>4|\mu|$. Using Lemma 10, we have that the sign of $Q(z^{\prime},w^{\prime})-Q(z,w)$ is the same as that of the inner product between the vector pointing from $(z,w)$ to $(z^{\prime},w^{\prime})$ and the normal vector $(-8\mu,Q(z,w))$ of the line $16\mu z-Q(z,w)^{2}+2Q(z,w)w=0$, which is given by
@@
$\{2\mu(z^{2}+\nu^{2}-2\nu)-(1-\nu)z(w-\sqrt{w^{2}-16\mu z})\leq 0\}$
$=$ $\{z^{2}\leq-\nu^{2}+2\nu\}\cup\left\{w<-\frac{\mu(z^{2}-2\nu+\nu^{2})}{z(\nu-1)}-\frac{4z^{2}(\nu-1)}{\nu^{2}-2\nu+z^{2}},z^{2}>-\nu^{2}+2\nu\right\},$
and
$\{2\mu(z^{2}+\nu^{2}-2\nu)-(1-\nu)z(w-\sqrt{w^{2}-16\mu z})=0\}$
$=$ $\left\{w=-\frac{\mu(z^{2}-2\nu+\nu^{2})}{z(\nu-1)}-\frac{4z^{2}(\nu-1)}{\nu^{2}-2\nu+z^{2}},z^{2}>-\nu^{2}+2\nu\right\},$
@@
Finally, when $\mu=0$, we have that $Q(z,w)=2w$. We have that
@@
Notice that $p(-\infty)=$-$\infty$ and $p(-1+\nu)=(\nu-1)^{2}(w_{0}-2(z_{0}+\mu))>0$. Hence, $p$ has at least one root in $(-\infty,-1+\nu)$. By Lemma 11, in particular, by (9), we have that, when viewing $F$ as a map on $\mathbb{R}^{2}$:
@@
We now show that $F$ maps $\partial A_{0}$ bijectively to $\partial\Omega$. Since $A_{0}\subset\mathrm{cl}(A_{0})$, we have that $F(A_{0})=\Omega^{o}\subset F(\mathrm{cl}(A_{0}))$. Since $\mathrm{cl}(A_{0})$ is compact, therefore, $F|_{\mathrm{cl}(A_{0})}\colon\mathrm{cl}(A_{0})\to\Omega$ is proper, and hence is closed (see, e.g., Theorem 4.95. Lee,, 2000). Therefore, $F(\mathrm{cl}(A_{0}))$ is a closed set that contains $\Omega^{o}$. Hence, $\mathrm{cl}(\Omega^{o})=\Omega\subset F(\mathrm{cl}(A_{0}))$. Since $F(A_{0})=\Omega^{o}$, we have $\partial\Omega\subset F(\partial A_{0})$, which means $F|_{\partial A_{0}}$ is onto $\partial\Omega$. We next show it is also injective. By Lemma 11, we know $F$ maps $\left\{z=\nu-1\right\}$ to $\left\{w=2(z+\mu)\right\}$ and maps $\left\{w=-2(z+\mu)\right\}$ to itself. When $z=\nu-1$, the $w$-update under $F$ is given by
@@
which is linear in $w$. Therefore, $F|_{z=1-\nu}$ must be an injection. When $w=-2(z+\mu)$, the $w$-update under $F$ is given by
@@
As a function $w$, $w^{\prime}$ have two critical points, $w=2(1-\mu-\nu)$ and $w=\frac{2}{3}(1-\mu-\nu)$. Notice that $z=\nu-1$ intersects $\partial\Omega$ at $(\nu-1,2(1-\mu-\nu))$. Then when $(z,w)\in\mathrm{cl}(A_{0})$, we have $w\geq 2(1-\mu-\nu)$. Since $1-\mu-\nu>0$, we have that the above $w^{\prime}$ is monotonic with $w$. Therefore, $F|_{\mathrm{cl}(A_{0})}$ is an injection. It follows that, $F|_{\partial A_{0}}$ is a bijection to $\partial\Omega$ and $F|_{\mathrm{cl}(A_{0})}$ is a bijection to $\Omega$. Since $F|_{\mathrm{cl}(A_{0})}$ is a closed map, its inverse is continuous. Therefore, $F|_{\mathrm{cl}(A_{0})}$ is a homeomorphism. The proof for $A_{2}$ is similar and thus is omitted.
@@
$6-4\nu-\frac{1-3z^{2}-2z\mu-2\nu+\nu^{2}}{1-\nu}>0$ (16)
$\Leftrightarrow$ $3z^{2}+2\mu z+(3\nu^{2}-8\nu+5)>0.$
For this parabola of $z$, we have
@@
Therefore, we have that (16) holds and that
@@
For this new parabola of $z$, we have its discriminant is negative if
@@ Appendix E Proofs for Section 3
In this section, we present the proofs of our main results. The key idea is to first analyze the quotient dynamical system introduced in Appendix D, and then translate the conclusions back to the original gradient descent system.
@@ E.1.1 Preliminary results
As discussed in Appendix D, the gradient descent system $\mathrm{GD}_{\eta}$ is semi-conjugate to the following system $F$:
@@
where $\mu=y\eta$ denote the parameter of the system and the state space of $F$ is
@@
In the following several results, we characterize the long term behavior of the orbits of $F$. We will use the terms trajectory and orbit interchangeably. We say an orbit $\{z_{k},w_{k}\}$ converges to a set $S$ if $d((z_{k},w_{k}),S)\to 0$, where $d(x,S)=\inf_{y\in S}d(x,y)$. Unless stated otherwise, we use $(z^{\prime},w^{\prime})$ to denote $F(z,w)$.
@@
  • If $w_{0}>\mu z_{0}+4$, the orbit either diverges, in the sense that $w_{k}\rightarrow\infty$, or converges to $\{z=0\}$ in finite steps.
@@
Consider the function $\Delta Q(z,w)=Q(F(z,w))-Q(z,w)$. Let $(z_{k},w_{k})=F^{k}(z_{0},w_{0})$ for $k\geq 1$. When $w_{0}<\mu z_{0}+4$, by Lemma 12, we have $Q(z_{k+1},w_{k+1})\leq Q(z_{k},w_{k})$ for all $k\geq 0$. Hence, the trajectory stays in the region $\left\{Q<8\right\}=\{w<\mu z+4\}$. Since $Q(z_{k},w_{k})$ is monotonic and is non-negative, it converges to some finite value and $\Delta Q(z_{k},w_{k})$ converges to zero. By Lemma 12, we have that
@@
If $(z_{k},w_{k})$ does not converges to $Z$, there exists $\varepsilon_{0}$ such that for any $K$ there exists $k>K$ such that $d((z_{k},w_{k}),Z)\geq\varepsilon_{0}$. Note that, the function $\Delta Q$, when restricted to the compact set $\left\{w\leq\mu z+4\colon\ d((z,w),Z)\geq\varepsilon_{0}\right\}$, is non-positive and continuous. Thus, it obtains its maximal value and the maximum is strictly negative. Hence, $d((z_{k},w_{k}),Z)\geq\varepsilon_{0}$ implies that $\Delta Q(z_{k},w_{k})<-\delta$ for some $\delta>0$, which contradicts the fact that $\Delta Q\to 0$. Therefore, $(z_{k},w_{k})$ converges to
@@
When $w_{0}>\mu z_{0}+4$, similarly, we have that $Q(z_{k+1},w_{k+1})\geq Q(z_{k},w_{k})$. Hence, $(z_{k},w_{k})$ stays in the region $\{w>\mu z+4\}$ for all $k\geq 0$. Hence, $Q(z_{k},w_{k})$ either diverges to infinity or converges to a finite value. If it diverges, $w_{k}$ must also diverge, since the function $Q$, when restricted to $\{w\leq\bar{w}\}$ for any fixed $\bar{w}$, is continuous and hence is upper bounded. If $Q(z_{k},w_{k})$ converges to some finite value, then $\Delta Q$ converges to zero and the trajectory must remain within the compact region $\{w\leq C\}$ for some $C>0$. Using arguments similar to those above, we have that $(z_{k},w_{k})$ must converge to $Z\cap\left\{w>\mu z+4\right\}=\left\{z=0\right\}$. Now assume the convergence is in infinite steps, i.e., $|z_{k}|\neq 0$ for all $k\in\mathbb{N}$. Then the sequence $|z_{k+1}/z_{k}|$ is well defined and converges to one. Notice that we have
$|\frac{z_{k+1}}{z_{k}}|=|z_{k}^{2}+\mu z_{k}-w_{k}+1|\geq\left||z_{k}^{2}+\mu z_{k}|-|w_{k}-1|\right|.$ (18)
Note as $z_{k}\to 0$, the above lower bound is dominated by $|w_{k}-1|$. Since $\{w>\mu z+4\}\cap\{z=0\}=\{(0,w)\colon w\geq 4\}$, $|\frac{z_{k+1}}{z_{k}}|$ is lower bounded by $1+\delta$ for some $\delta>0$. This contradicts the fact that $|z_{k+1}/z_{k}|$ converges to one. Hence, the convergence must occur within finite steps.
Finally, the result for the case $w_{0}=\mu z_{0}+4$ directly comes from Lemma 12. This completes the proof. ∎
@@
Consider $|\mu|<1$. By Proposition 14, when $Q(z,w)>8$, initializations either converge to $\{z=0\}$ in finite steps or diverge. Notice that converging within finite steps means that the initialization lies in the set